OpenStudy (anonymous):

giving medal!! help me!!???:O The statement e + (f + g) = ( f + g) + e is an example of which property of addition? a. commutative c. identity b. associative d. substitution

OpenStudy (mathstudent55):

What has changed, the grouping or the order?

hartnn (hartnn):

commutative : m*n=n*m identity : a*1 =a associative : m+ (n+p) = (m+n) + p now choose :)

OpenStudy (anonymous):

commutative?

OpenStudy (mathstudent55):

Change of order ---> commutative Change of grouping ---> associative

OpenStudy (mathstudent55):

correct

OpenStudy (mathstudent55):

It IS commutative. The grouping has not changed. It is the order that changed. Look at what is inside the parentheses on both sides of the equal sign.

hartnn (hartnn):

yeah, sorry, its commutative :)

OpenStudy (mathstudent55):

The use of parentheses in this question is to diguise the commutative property into looking like the associative property.

OpenStudy (mathstudent55):

Thanks, @hartnn